OpenStudy (anonymous):

Please show work if possible :D Thank you 6 1/3 divided by * (-3 8/9) =???

OpenStudy (anonymous):

Can any one help me???

OpenStudy (anonymous):

Hold on, I'll help.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

hello...

OpenStudy (anonymous):

Oh, I'm sorry I thought it posted! 6 1/3 as an improper is 19/3. -3 8/9 as an improper is -216/9. Keep it, Change it, Flip it. -171/648.

OpenStudy (anonymous):

Its OK Hehe :D

OpenStudy (anonymous):

is that symplyfied? :D

OpenStudy (anonymous):

13/18 * 20/27 = ???Please show work...can you help me please :(

OpenStudy (anonymous):

-57/216 would be simplified.

OpenStudy (anonymous):

Multiply across. 13*20=260. 18/27=486. 260/486. 130/243. :)